# Sensing with near-infrared laser trapped fluorescent nanodiamonds

### v1

@@ III.6 Temperature Simulation
  • Steady-State Heat Transfer Coefficient: The convective heat transfer coefficient $h$ is considered constant and based on the thermal conductivity of the medium.
@@
Consistency with Experimental Observations
The simulation results align with experimental observations where pulsed NIR illumination causes negligible thermal effects in aqueous environments. This supports the conclusion that charge state dynamics are the primary mechanism affecting NV center behavior under short NIR pulses in water.
@@
In experiments conducted in air, temperature increases contribute to shifts in ODMR frequency and should be accounted for. These shifts correspond to a 2 MHz shift in the ODMR resonance frequency which is what we experimentally observe.
IV Discussion: Effects of NIR Light on FNDs
Our experimental and simulation results reveal a complex interplay between thermal dynamics and charge state transitions in NV within FNDs under NIR illumination. The observed effects include a decrease in total fluorescence intensity, reduction in spin-lattice relaxation time, shift in the ODMR resonance frequency, decrease in ODMR contrast, and an increase in the NV0/NV- fluorescence ratio.
IV.1 Thermal Effects Induced by NIR Illumination
One primary mechanism contributing to these observations is the localized heating induced by the NIR laser. The shift in the ODMR resonance frequency with increasing NIR power, as shown in Figures 6, 7 and 10, and 11a), suggests an increase in temperature within the FNDs. This is consistent with the known temperature dependence of the zero-field splitting parameter $D$ of the NV center, which decreases with increasing temperature at a rate of approximately $-74$ kHz/K [33, 34].
However, the extent of the thermal effect is significantly influenced by the surrounding medium’s thermal properties. Media with higher thermal conductivity and heat capacity, such as water, dissipate heat more efficiently than air, resulting in smaller temperature increases in the FNDs. This is supported by our temperature simulations, which show minimal temperature rise in water compared to air during NIR illumination. Consequently, FNDs trapped in water exhibit much smaller ODMR frequency shifts and negligible changes in $T_{1}$ times with increasing NIR power, as shown in Figures 7, 10 and 11.
IV.2 Charge State Dynamics and Fluorescence Changes
In addition to thermal effects, the NIR laser alters the charge state dynamics of the NV centers. Our simulations of charge state dynamics indicate that at low NIR power levels, there is a slight increase in the NV- population due to enhanced recombination processes facilitated by NIR photons. However, as the NIR power increases beyond a threshold (approximately 10 mW with our parameters), the NV- population begins to decrease while the NV0 population increases. This inversion is attributed to the saturation of recombination processes and increased ionization rates at higher NIR powers.
These changes in charge state populations directly affect the fluorescence properties of the FNDs. The total fluorescence intensity decreases with increasing NIR power, as shown in Figure 13. This decrease can be explained by the reduced NV- population and the lower fluorescence quantum yield of NV0 centers [35]. The increase in the NV0/NV- fluorescence ratio with higher NIR power, depicted in Figures 10 and 11, further confirms the shift in charge state distribution.
Additionally, the increased NV0 background fluorescence dilutes the ODMR contrast since NV0 centers do not exhibit spin-dependent fluorescence variations. This is reflected in the decreased ODMR contrast observed in our measurements (Figures 6-7). Charge-state fluctuations also introduce additional pathways for spin relaxation, contributing to the reduction in $T_{1}$ times [20, 36].
IV.3 Interplay Between Thermal Effects and Charge State Dynamics
Although both thermal effects and charge state dynamics significantly impact the properties of the NV centers under NIR illumination, their relative contributions depend on the environmental conditions. In FNDs dried on glass substrates, the pronounced ODMR frequency shifts suggest that thermal effects dominate due to inefficient heat dissipation in air.
In contrast, for FNDs trapped in water, the minimal temperature increase reduces thermal effects, making charge-state dynamics more prominent in influencing the NV centers’ behavior. The increase in the NV0/NV- ratio and the decrease of T1-relaxometry times with NIR power without significant ODMR frequency shifts (Figure 10) indicates that charge state transitions are the primary mechanism affecting the fluorescence and spin properties in aqueous environments.
IV.4 Alternative Explanations and Considerations
Although thermal effects and charge state dynamics provide a comprehensive explanation for our observations, alternative mechanisms should be considered. Multi-photon absorption processes could directly excite the NV centers without significant heating. However, the relationship between ODMR frequency shift and NIR power (Figures 7 and 10) supports the thermal interpretation in our experiments.
Moreover, the significant differences in the effects of NIR irradiation between FNDs in air and water emphasize the critical role of the surrounding medium. The statistical analysis using ANOVA (Table 1) confirms that the differences observed in FNDs dried on glass are statistically significant, whereas those in FNDs trapped in water are not, reinforcing the influence of environmental thermal properties.
V Sensing Capabilities of Trapped FNDs
@@
As a final step in the comprehensive investigations, this section explores the potential of optically trapped FNDs with NV centers for sensing applications. Experiments focused on three key parameters: Gd3+ concentration, pH, and temperature. Measurements included ODMR, T1 relaxation, and optical spectrum analysis under variation of the three parameters mentioned above.
@@
Figure 14: Variation in three sensing parameters when subject to controlled variation of paramagnetic species, pH, and temperature. a) ODMR resonance frequency shift VS Gd3+ concentration. b) ODMR resonance frequency shift VS pH value. c) ODMR resonance frequency shift VS Temperature increase. d) T1-relaxometry time VS Gd3+ concentration. e) T1-relaxometry time VS pH value. f) T1-relaxometry time VS Temperature increase. g) NV0/NV- fluorescence intensity ratio VS Gd3+ concentration. h) NV0/NV- fluorescence intensity ratio VS pH value. i) NV0/NV- fluorescence intensity ratio VS Temperature increase. The boxplot represent results for 10 FNDs that were trapped for each condition with 60 mW 1064 nm CW NIR laser. In each case, results for Gd3+ 0M, pH 13 and 0K trapped were used as reference values for the individual FNDs. The red squares emphasize the optimal measurement method for each sensing parameter. The ODMR resonance frequency value is given relative to the expected 2.87 GHz.
V.1 pH Sensing
@@
Figure 15: NV0/NV- fluorescence intensity ratio during pH value changes. Measurement of the same FND while changing from pH13 to pH1 by adding HCl solution (Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
  • T1 Relaxation: A decrease in T1 values was observed as pH increased, indicating a dependence on surface protonation dynamics (Fig. 14.h). This is consistent with reports that high pH conditions stabilize spin relaxation dynamics [1].
  • Optical Spectrum: A decrease in the NV0/NV- fluorescence intensity ratio with increasing pH (Fig. 14.e and  15) suggests that alkaline conditions favor the NV- charge state, possibly due to deprotonation of surface functional groups, which enhances negative surface potential [37]. The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser shows a clear advantage of using FNDs in suspension for sensing.
@@
The NV0/NV- fluorescence intensity ratio’s dependence on pH underscores the importance of surface chemistry in modulating charge state dynamics, especially with FNDs in suspension.
V.2 Paramagnetic Species Sensing
@@
Figure 16: NV0/NV- fluorescence intensity ratio during gadolinium concentration changes. Measurement of the same FND while changing gadolinium concentration from 0M to 0.1M by adding a gadolinium solution (0.45 M in acetate) (Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
  • T1 Relaxation: A significant decrease in T1 values was observed with increasing Gd3+ concentration (Fig. 14.g). This result aligns with the quenching effect of paramagnetic ions, which enhance spin-lattice relaxation by introducing magnetic noise to the local environment [38].
  • Optical Spectrum: A decrease in the NV0/NV- fluorescence intensity ratio was observed (Fig. 14.d and  16). This can be attributed to charge state dynamics, where Gd3+ ions interact with the FND surface, altering the local electrochemical potential and promoting ionization of NV- to NV0 [39]. The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser shows a clear advantage of using FNDs in suspension for sensing.
@@
These findings demonstrate the potential of NV centers for detecting paramagnetic species like Gd3+. The sensitivity saturates at higher concentrations due to the finite interaction volume and saturation of surface sites for Gd3+ binding. The T1-relaxometry provides a robust method for quantifying Gd3+ in biological and environmental contexts. However, the effect on charge state dynamics also imply that conclusions from T1-relaxometry in otherwise uncharacterized environments should be made with caution, and at best complemented with fluorescence spectrometry. For completeness, we observe that by adding Gd3+ to a concentration of 0.1 M, the pH value in the solution increases slightly (by of order 1-2 pH units) which is consistent with our previous observations on the effect of pH on the fluorescence of FNDs..
V.3 Temperature Sensing
@@
[…]
@@
  • T1 Relaxation: No significant changes were detected in T1 values, indicating that spin-lattice relaxation is relatively insensitive to temperature variations within the tested range (Fig. 14.i).
@@
  • Optical Spectrum: No significant changes in the NV0/NV- fluorescence intensity ratio were observed with temperature, suggesting that charge state dynamics are minimally affected by thermal fluctuations (Fig. 14.f).
@@
ODMR’s temperature sensitivity highlights its utility for high-resolution nanoscale thermometry. The absence of significant T1 or optical spectral changes simplifies calibration, making ODMR a standalone method for temperature measurements
@@ VI Conclusion
This work investigates the effects of near-infrared (NIR) laser irradiation on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers, offering a comprehensive understanding of the interplay between thermal effects, charge state dynamics, and their combined influence on sensing capabilities. By employing an integrated experimental approach combining photoluminescence spectroscopy, optically detected magnetic resonance (ODMR), and T1 relaxometry, this study elucidates how varying NIR laser powers impact the fluorescence, spin properties, and stability of NV centers in diverse environmental conditions.
A key finding is the dual role of NIR irradiation: in dry environments, thermal effects dominate, resulting in significant shifts in ODMR frequencies and reductions in T1 relaxation times due to localized heating. In contrast, for FNDs that are fully submerged in aqueous environments, the surrounding water appears to mitigate these thermal effects through efficient heat dissipation, allowing charge state dynamics to take precedence. The increased NV-, NV0 fluorescence intensity ratio and its implications for sensing applications underscore the critical role of the surrounding medium in tailoring NV center behavior. The study also demonstrates the sensitivity of optically trapped FNDs to pH variations, temperature fluctuations, and paramagnetic species, establishing these nanodiamonds as versatile tools for nanoscale sensing.
The results have profound implications for the development of robust NV-based sensing protocols. The findings indicate that optimizing NIR laser parameters, including power levels and pulsing sequences, can significantly enhance sensor performance while minimizing adverse effects such as fluorescence quenching and spin relaxation. Additionally, the demonstrated ability to detect and differentiate environmental factors such as pH and gadolinium ions highlights the potential of NV centers for applications in biological sensing and environmental monitoring.
@@
Theoretical modeling, while providing valuable insights into charge state dynamics and thermal effects, must be expanded to include the complexities of real-world applications. Future work could explore the integration of complementary sensing techniques, such as fluorescence lifetime imaging or advanced magnetic resonance methods, to provide a more holistic understanding of NV center behavior under varying conditions. Furthermore, efforts to miniaturize and automate the experimental setup, enabling in vivo applications, would be a critical step toward translating this research into practical diagnostic tools.
In conclusion, this study lays a strong foundation for the continued development of NV-based sensing technologies, bridging fundamental insights with application-driven innovations. By addressing the remaining uncertainties and refining experimental protocols, FNDs with NV centers have the potential to become transformative tools in nanoscale sensing, advancing fields ranging from cellular biology to nanoscale physics.
Acknowledgements.
This work was supported by the Independent Research Fund Denmark (grant no. 0135-00142B) and the Novo Nordisk Foundation (grant no. NNF20OC0061673). We acknowledge discussions with Maabur Sow and Fedor Jelezko, supported by ERC through HyperQ Project (SyG 856432). We further acknowledge insights provided through the student project work of Kristian Lambertsen and Nicolai Kongstad, including the sketch of the experimental setup, Fig. 2.

### v2

@@ III.5 Temperature Simulation
  • Steady-State Heat Transfer Coefficient: The convective heat transfer coefficient is considered constant and based on the thermal conductivity of the medium.
@@
The simulation results align with experimental observations where pulsed NIR illumination causes negligible thermal effects in aqueous environments, cf. Fig. 7. In experiments conducted in air, temperature increases contribute to shifts in ODMR frequency and should be accounted for. These shifts correspond to a 2 MHz shift in the ODMR resonance frequency which is what we experimentally observe.
A direct comparison of Fig. 6c)–d) with Fig. 6a)–b) shows that the environment (air versus water) is the controlling parameter. In air, the ODMR resonance shifts by $5.8\pm 0.5$ MHz and $T_{\rm relax}$ falls by $(44\pm 6)\%$ when the CW NIR power is raised from 0 to 60 mW, whereas in water the corresponding changes are only $1.6\pm 0.3$ MHz and $(2\pm 5)\%$, respectively. By contrast, the NV0/NV- fluorescence intensity ratio $\sigma_{0-}$ responds similarly to NIR power in both media (Fig. 6a,c; projection on the ordinate axis), confirming that charge-state kinetics are governed by photon flux rather than heat removal.
For completeness we note that NV- centres possess a weak intersystem transition between the metastable singlet states $^{1}E\,(p_{6})\!\rightarrow\!{}^{1}A_{1}\,(p_{5})$ with a zero-phonon line at $\lambda_{\mathrm{ZPL}}=1042$ nm and a room-temperature homogeneous width of only a few GHz ($\lesssim 10^{-4}$ nm) [Acosta2010, Rogers2015]. Our trapping beam at 1064 nm is therefore red-detuned by $\Delta\lambda=22$ nm, corresponding to $\Delta\nu\simeq 6$ THz ($\simeq$  25 meV), i.e. $\sim 10^{3}$ times the natural linewidth and well outside the vibrational sideband of the singlet manifold. Values for the measured absorption cross-sections decrease from $\sigma\!\approx\!2\times 10^{-19}$ cm2 on resonance to below $10^{-21}$ cm2 at this detuning [Acosta2010], so direct optical pumping of the singlet transition by the 1064 nm light is negligible. The effects reported in this work (PL quenching, ODMR shifts, $T_{\rm relax}$ variations) therefore arise from photothermal heating and photo-ionisation dynamics, not from resonant excitation of the 1042 nm line.
IV Sensing Capabilities of Trapped FNDs
@@
As a final step in the comprehensive investigations, this section explores the potential of optically trapped FNDs with NV centers for sensing applications. Experiments focused on three key parameters: Gd3+ concentration, pH, and temperature. Measurements included ODMR, fluorescence relaxation, and optical spectrum analysis under variation of the three parameters mentioned above. Results are summarized in Fig. 10.
@@
Figure 10: Variation in three sensing parameters when subject to controlled variation of paramagnetic species, pH, and temperature. a) ODMR resonance frequency shift VS Gd3+ concentration. b) ODMR resonance frequency shift VS pH value. c) ODMR resonance frequency shift (relative to the expected 2.87 GHz) VS Temperature increase (relative to 37 C). d) Fluorescence relaxometry time $T_{\rm relax}$ VS Gd3+ concentration. e) $T_{\rm relax}$ VS pH value. f) $T_{\rm relax}$ VS Temperature increase. g) $\sigma_{0-}$ fluorescence intensity ratio VS Gd3+ concentration. h) $\sigma_{0-}$ fluorescence intensity ratio VS pH value. i) $\sigma_{0-}$ fluorescence intensity ratio VS Temperature increase. The boxplot represent results for 10 FNDs that were trapped for each condition with 60 mW 1064 nm CW NIR laser. In each case, results for 0 M Gd3+, pH 13 and 0 K trapped were used as reference values for the individual FNDs. The red squares emphasize the optimal measurement method for each sensing parameter.
IV.1 pH Sensing
@@
Figure 11: $\sigma_{0-}$ fluorescence intensity ratio during pH value changes. Measurement of the same FND while changing from pH13 to pH1 by adding HCl solution. Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
  • Fluorescence relaxation: A decrease in $T_{\rm relax}$ values was observed as pH increased, indicating a dependence on surface protonation dynamics (Fig. 10.h). This is consistent with reports that high pH conditions stabilize spin relaxation dynamics [Fujisaku2019, Rendler2017].
  • Optical Spectrum: A decrease in the $\sigma_{0-}$ ratio with increasing pH (Fig. 10.e and Fig. 11) suggests that alkaline conditions favor the NV- charge state. It may be attributed to the deprotonation of surface functional groups such as carboxyl (–COOH) and hydroxyl (–OH) groups present on the nanodiamond surface. Deprotonation of these groups under alkaline conditions generates negatively charged surface sites, thereby enhancing electron donation to nearby NV centers and favoring the stabilization of the NV- charge state [karaveli2016]. The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser (Fig. 11) shows a clear advantage of using FNDs in suspension for sensing.
@@
The $\sigma_{0-}$ ratio’s dependence on pH underscores the importance of surface chemistry in modulating charge state dynamics, especially with FNDs in suspension.
To further explore the dependence of the NV charge state on pH, Fig. 11 presents the evolution of the $\sigma_{0-}$ ratio when varying pH from 13 to 1 for the same FND, both under trapping conditions and without NIR laser exposure. The trapped FNDs in suspension exhibit a significantly stronger modulation of $\sigma_{0-}$ compared to the same particle measured on glass. This enhanced responsiveness highlights the improved surface accessibility and chemical exchange in solution, reinforcing the utility of optically trapped FNDs for pH nanosensing. The pronounced increase in NV center fraction under alkaline conditions (lower $\sigma_{0-}$) confirms the surface deprotonation hypothesis, while acidic conditions suppress this stabilization. Overall, these results validate that $\sigma_{0-}$ serves as a reliable and sensitive optical proxy for probing local pH variations in biological or environmental settings.
IV.2 Paramagnetic Species Sensing
@@
Figure 12: $\sigma_{0-}$ fluorescence intensity ratio during gadolinium concentration changes. Measurement of the same FND while changing gadolinium concentration from 0M to 0.1M by adding a gadolinium solution (0.45 M in acetate). Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
  • Fluorescence relaxation: A significant decrease in $T_{\rm relax}$ values was observed with increasing Gd3+ concentration (Fig. 10.g). This result aligns with the quenching effect of paramagnetic ions, which enhance spin-lattice relaxation by introducing magnetic noise to the local environment [Martinez2020, Rendler2017].
  • Optical Spectrum: A decrease in $\sigma_{0-}$ was observed (Fig. 10.d and Fig. 12). This can be attributed to charge state dynamics, where Gd3+ ions interact with the FND surface, altering the local electrochemical potential and promoting ionization of NV- to NV0 [Rondin2010]. The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser (fig. 12) shows a clear advantage of using FNDs in suspension for sensing.
@@
These findings demonstrate the potential of NV centers for detecting paramagnetic species like Gd3+. The sensitivity saturates at higher concentrations due to the finite interaction volume and saturation of surface sites for Gd3+ binding. The fluorescence relaxometry provides a robust method for quantifying Gd3+ in biological and environmental contexts. However, the effect on charge state dynamics also imply that conclusions from fluorescence relaxometry in otherwise uncharacterized environments should be made with caution, and at best complemented with fluorescence spectrometry. For completeness, we observe that by adding Gd3+ to a concentration of 0.1 M, the pH value in the solution increases slightly (by of order 1-2 pH units) which is consistent with our previous observations on the effect of pH on the fluorescence of FNDs.
IV.3 Temperature Sensing
@@
  • ODMR: A temperature-dependent shift in the resonance frequency was observed, with lower frequencies at higher temperatures (Fig. 10.c). This shift is attributed to the well-known temperature dependence of the NV center’s zero-field splitting parameter $D$, which decreases at a rate of approximately $74~\text{kHz/K}$ [Acosta2010].
@@
  • Fluorescence relaxation: No significant changes were detected in $T_{\rm relax}$ values, indicating that spin-lattice relaxation is relatively insensitive to temperature variations within the tested range (Fig. 10.i).
@@
  • Optical Spectrum: No significant changes in $\sigma_{0-}$ were observed with temperature, suggesting that charge state dynamics are minimally affected by thermal fluctuations (Fig. 10.f).
@@
ODMR’s temperature sensitivity highlights its utility for high-resolution nanoscale thermometry. The absence of significant $T_{\rm relax}$ or optical spectral changes simplifies calibration, making ODMR a standalone method for temperature measurements
V Discussion
The shift in the ODMR resonance frequency with increasing NIR power, as shown in Fig. 4, Fig. 6, and Fig. 7a), suggests an increase in temperature within the FNDs. This is consistent with the known temperature dependence of the zero-field splitting parameter $D$ of the NV center, which decreases with increasing temperature at a rate of approximately 74-74 kHz/K [Acosta2010, Toyli2012]. However, the extent of the thermal effect is significantly influenced by the surrounding medium’s properties. Media with higher thermal conductivity and heat capacity, such as water, dissipate heat more efficiently than air, resulting in smaller temperature increases in the FNDs. This is supported by our temperature simulations, which show minimal temperature rise in water compared to air during NIR illumination. Consequently, FNDs trapped in water exhibit much smaller ODMR frequency shifts with increasing NIR power, as shown in Fig. 6 and Fig. 7.
In addition to thermal effects, the NIR laser alters the charge-state dynamics of the NV centers. Our simulations of charge-state dynamics indicate that at low NIR power levels, there is a slight increase in the NV- population due to enhanced recombination processes facilitated by NIR photons. However, as the NIR power increases beyond a threshold (approximately 10 mW with our parameters), the NV- population begins to decrease while the NV0 population increases. This inversion is attributed to the saturation of recombination processes and increased ionization rates at higher NIR powers.
The total fluorescence intensity decreases with increasing NIR power, as shown in Fig. 9. This decrease can be explained by the reduced NV- population and the lower fluorescence quantum yield of NV0 centers [Chapman2010]. The increase in $\sigma_{0-}$ with higher NIR power, depicted in Fig. 6 and Fig. 7, further confirms the shift in charge state distribution. Additionally, the increased NV0 background fluorescence dilutes the ODMR contrast since NV0 centers do not exhibit spin-dependent fluorescence variations. This is reflected in the decreased ODMR contrast observed in our measurements (Fig. 4). Charge-state fluctuations also introduce additional pathways for spin relaxation, contributing to the reduction in $T_{\rm relax}$ times [Aslam2013, Goudarzi2019].
Our results indicate that although both thermal effects and charge-state dynamics significantly impact the properties of the NV centers under NIR illumination, their relative contributions depend on the environmental conditions. In FNDs dried on glass substrates, the pronounced ODMR frequency shifts suggest that thermal effects dominate due to inefficient heat dissipation in air. In contrast, for FNDs trapped in water, the minimal temperature increase reduces thermal effects, making charge-state dynamics more prominent in influencing the NV centers’ behavior. The increase in the $\sigma_{0-}$ ratio and the decrease of fluorescence relaxometry times with NIR power without significant ODMR frequency shifts (Fig. 6) indicate that charge-state transitions are the primary mechanism affecting the fluorescence and spin properties in aqueous environments.
While thermal effects and charge-state dynamics provide a comprehensive explanation for our observations, alternative mechanisms should be considered. Multi-photon absorption processes could directly excite the NV centers without significant heating, yet, the relationship between ODMR frequency shift and NIR power (Fig. 6) is contrary to this explanation. Moreover, the significant differences in the effects of NIR irradiation between FNDs in air and water emphasize the critical role of the surrounding medium. The statistical analysis using ANOVA (Table 1) confirms that the differences observed in FNDs dried on glass are statistically significant, while those in FNDs trapped in water are not.
In our opinion, the data presented in this work establish a clear hierarchy of influences: (i) the thermal conductivity and heat capacity of the surrounding medium dictate the magnitude of the NIR-induced temperature rise and therefore the thermally induced shifts in ODMR frequency and the shortening of $T_{\rm relax}$; (ii) the NIR photon flux sets the rate of photo-ionisation and recombination, which in turn controls the NV0/NV- balance and the fluorescence contrast. Optical trapping itself plays a mainly mechanical role-holding a single FND at a fixed position—while potential effects on spin and charge-state dynamics depend on the wavelength of the trapping laser.
This distinction is important for two reasons. First, it tells experimentalists which knob to turn: to suppress thermal artefacts one changes the medium (e.g. buffer composition, flow, heat sinking) rather than the trap geometry and wavelength. Second, it generalises our findings to intracellular sensing, where the cytoplasm’s thermal properties resemble those of water: the modest $\lvert\Delta f\rvert<2$ MHz and unchanged $T_{\rm relax}$ observed in vitro set an upper bound on in-cell NIR heating, validating the use of trapped FNDs for quantitative biosensing.
The results presented in Section IV confirm that optical trapping of FNDs provide a valuable add-on that allows to enhance the application potential of biosensing with NV-centers in nanodiamonds. In particular, Fig. 10a-c) demonstrates that the frequency shift in ODMR is a good stand-alone measure for a temperature shift also for optically trapped FNDs, while parts f) and i) illustrate that changes in the surrounding temperature within a range that is relevant for biological systems does not appear to alter the readings provided by the fluorescence intensity ratio $\sigma_{0-}$ or the relaxation time, $T_{\rm relax}$.
@@ VI Conclusion
This work investigates the effects of near-infrared (NIR) laser irradiation on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers, offering a comprehensive understanding of the interplay between thermal effects, charge-state dynamics, and their combined influence on sensing capabilities. By employing an integrated experimental approach combining photoluminescence spectroscopy, optically detected magnetic resonance (ODMR), and fluorescence relaxometry, this study elucidates how varying NIR laser powers impact the fluorescence, spin properties, and stability of NV centers in diverse environmental conditions.
A key finding is the dual role of NIR irradiation: in dry environments, thermal effects dominate, resulting in significant shifts in ODMR frequencies and reductions in relaxation times due to localized heating. In contrast, for FNDs that are fully submerged in aqueous environments, the surrounding water appears to mitigate these thermal effects through efficient heat dissipation, allowing charge state dynamics to take precedence. The increased $\sigma_{0-}$ ratio and its implications for sensing applications underscore the critical role of the surrounding medium in tailoring NV center behavior. The study also demonstrates the sensitivity of optically trapped FNDs to pH variations, temperature fluctuations, and paramagnetic species, establishing these nanodiamonds as versatile tools for nanoscale sensing.
The findings indicate that optimizing NIR laser parameters, including power levels and pulsing sequences, can significantly enhance sensor performance while minimizing adverse effects such as fluorescence quenching and spin relaxation. Additionally, the demonstrated ability to detect and differentiate environmental factors such as pH and gadolinium ions highlights the potential of NV centers for applications in biological sensing and environmental monitoring.
@@
Theoretical modeling, while providing valuable insights into charge state dynamics and thermal effects, must be expanded to include the complexities of real-world applications. Future work could explore the integration of complementary sensing techniques, such as fluorescence lifetime imaging or advanced magnetic resonance methods, to provide a a real-time sensing protocol inside cells by using optically trap FNDs in the cellular medium. Furthermore, efforts to miniaturize and automate the experimental setup, enabling in vivo applications, would be a critical step in translating this research into practical diagnostic tools.
acknowledgments
This work was supported by the Independent Research Fund Denmark (grant no. 0135-00142B) and the Novo Nordisk Foundation (grant no. NNF20OC0061673). We acknowledge discussions with Maabur Sow and Fedor Jelezko, supported by ERC through HyperQ Project (SyG 856432). We further acknowledge insights provided through the student project work of Kristian Lambertsen and Nicolai Kongstad, as well as Suyash Amzare, including an original sketch of the experimental setup providing the basis for Fig. 2.